# General Preference Modeling with Preference Representations for Aligning Language Models

### v1

@@ Abstract
Modeling human preferences is crucial for aligning foundation models with human values. Traditional reward modeling methods, such as the Bradley-Terry (BT) reward model, fall short in expressiveness, particularly in addressing intransitive preferences. Although supervised pair preference models (PairPM) can express general preferences, their implementation is highly ad-hoc and cannot guarantee a consistent preference probability of compared pairs. Additionally, they impose high computational costs due to their quadratic query complexity when comparing multiple responses. In this paper, we introduce preference representation learning, an approach that embeds responses into a latent space to capture intricate preference structures efficiently, achieving linear query complexity. Additionally, we propose preference score-based General Preference Optimization (GPO), which generalizes reward-based reinforcement learning from human feedback. Experimental results show that our General Preference representation model (GPM) outperforms the BT reward model on the RewardBench benchmark with a margin of up to 5.6% and effectively models cyclic preferences where any BT reward model behaves like a random guess. Furthermore, evaluations on downstream tasks such as AlpacaEval2.0 and MT-Bench, following the language model post-training with GPO and our general preference model, reveal substantial performance improvements with margins up to 9.3%. These findings indicate that our method may enhance the alignment of foundation models with nuanced human values. The code is available at https://github.com/general-preference/general-preference-model.
@@ 1 Introduction
Modeling human preferences is a cornerstone in developing foundation models that interact seamlessly with users. In natural language modeling and reinforcement learning, aligning models with human intent and values has led to significant advancements, including improved text generation and enhanced decision-making policies (Ouyang et al., 2022; Christiano et al., 2017). Traditional approaches often rely on reward modeling, wherein a reward function is learned to guide the optimization of policies. While effective in certain contexts, these methods face challenges in expressiveness and computational efficiency, particularly when addressing complex or intransitive human preferences (Tversky, 1969; Munos et al., 2023).
Figure 1: Intransitiveness in real-world preferences.
@@
(b) PairPM
(c) General Preference representation model (GPM)
Figure 2: Illustration of (a) Bradley Terry (BT) reward model, (b) supervised pair preference model (PairPM) (Jiang et al., 2023; Dong et al., 2024), and (c) our General Preference representation model (GPM).
@@
On the other hand, supervised (sequential-classification) pair preference models (PairPM) (Jiang et al., 2023; Dong et al., 2024) that predict the preference given a concatenation of the two responses can express complex and intransitive (cyclic) structures. But to fully capture the preference relations among $K$ responses, it requires evaluating $\mathcal{O}(K^{2})$ pairwise preferences between all $K$ candidate responses (Munos et al., 2023; Wu et al., 2024b). This quadratic scaling hinders them for applications with larger response sets especially in test-time scaling for reasoning tasks using verifiers and ranking models (Snell et al., 2024; Wu et al., 2024a).
Aside from computational inefficiency, supervised preference models also exhibit asymmetric preference behaviors related to positions. Also, the model’s design choice can be highly ad-hoc, varying among different templates and different linear heads.
@@
In this paper, we answer this question affirmatively by proposing preference representation learning, which bridges the gap between expressiveness and efficiency in general preference modeling. Our method embeds responses into a multi-dimensional latent space that captures the complex preference structure beyond transitive relations while allowing for efficient querying of preferences. Notably, our approach achieves a computational complexity of $\mathcal{O}(K)$, matching the efficiency of the BT model but with enhanced expressiveness.
@@
  • We introduce preference representation learning for general preference modeling, enabling both efficient and expressive representation of human preferences. Our approach generalizes the Bradley-Terry (BT) reward model by embedding responses into a latent space, capturing complex preference structures, including intransitive preferences. Notably, our General Preference representation model (GPM) achieves a query complexity of $\mathcal{O}(K)$ for evaluating preferences among $K$ responses, a significant improvement over the $\mathcal{O}(K^{2})$ complexity of traditional supervised preference models that rely on pairwise inputs (see Section 4).
  • We demonstrate the effectiveness of GPM across a variety of tasks, including CyclicPreference (ours) and the renowned RewardBench (Lambert et al., 2024). Specifically, GPM models intransitive (e.g., cyclic) preferences with 100% accuracy, whereas the BT reward model performs like random guessing (see Section 6.1). Additionally, GPM outperforms the BT model on RewardBench with performance margins of up to 5.6% (see Section 6.2).
  • For language model alignment, we propose General Preference Optimization (GPO), which leverages the preference scores provided by GPM. The general preference score can also be integrated as a preference signal into a wide range of RLHF and preference optimization methods, such as (iterative) DPO (Rafailov et al., 2024), SPPO (Wu et al., 2024b), and PPO-based methods (Ouyang et al., 2022). Experimental results on AlpacaEval2.0 and MT-Bench reveal that our approach leads to substantial improvements over reward-based methods for language model alignment (see Section 6.3).
@@ 2 Related Work
Reward-Based Reinforcement Learning from Human Feedback (RLHF). The earlier approaches to modeling human preference for language model alignment usually learn a reward model from a preference dataset. The human preference is assumed to follow the Bradley-Terry (BT) model (Bradley & Terry, 1952) or the Thurstone model (Thurstone, 2017). LLM policies then are fine-tuned to maximize these scalar reward signals for better alignment (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022). Later, the direct preference optimization (DPO) methods are proposed by Rafailov et al. (2024) to only implicitly learn a reward model represented by an LLM. The human preference is still assumed to follow the Bradley-Terry model. However, the reliance on scalar rewards imposes a total ordering on preferences, which may not reflect the intransitive or stochastic nature of human judgments (Tversky, 1969; Agranov & Ortoleva, 2017).
@@
Intransitivity in Game Theory.
The symmetric zero-sum game and its intransitivity have also been frequently studied in the context of game theory. Balduzzi et al. (2018) was motivated by evaluation among different agents, showing that any symmetric zero-sum game can be decomposed into a “transitive” game and a “cyclic” game, and proposed Nash averaging for better agent/task evaluation. Balduzzi et al. (2019) generalized the results from matrix games to functional-form games and propose new algorithms to construct diverse populations of effective agents. Czarnecki et al. (2020) investigated the geometrical properties of real-world games (e.g., Tic-Tac-Toe, Go, StarCraft II) and proposed that real-world games have a “spinning top” geometry, with a strong transitive dimension and gradually diminishing non-transitive cyclic dimensions. Very recently, Bertrand et al. (2023) examined the limitations of the Elo rating system and proposed an alternative “disc decomposition” method that can better handle both transitive and cyclic game dynamics.
Representation Learning and Embedding.
Representation learning and embedding techniques have successfully captured relational structures across various domains (Mikolov et al., 2013; Chen et al., 2020; Radford et al., 2021), yet their application in preference modeling and RLHF remains limited. Our work introduces preference representation learning, an approach that enhances expressiveness while maintaining computational efficiency, bridging the gap left by traditional approaches.
@@ 3 Background
In this section, we present preliminaries on reward modeling, preference modeling, and reinforcement learning from human feedback (RLHF) for language model alignment. We consider an autoregressive language model that generates responses to the given prompts. Let $\mathbf{x}=[x_{1},x_{2},\ldots]$ denote a prompt, a sequence of tokens. The language model $\pi$ generates a response $\mathbf{y}=[y_{1},y_{2},\ldots,y_{N}]$ based on the conditional probability distribution: $\pi(\mathbf{y}\mid\mathbf{x})=\prod_{i=1}^{N}\pi\left(y_{i}\mid\mathbf{x},\mathbf{y}_{<i}\right)$, where $\mathbf{y}_{<i}$ represents the sequence of tokens generated before position $i$. In this paper, we assume a general-preference oracle. Given two responses $\mathbf{y}$ and $\mathbf{y}^{\prime}$ to the same prompt $\mathbf{x}$, the oracle provides the feedback indicating which response is preferred.
@@ 3.1 Reward-Based Reinforcement Learning from Human Feedback
The most prevalent approach to aligning language models with human preferences is to consider a scalar reward function $r(\mathbf{y};\mathbf{x})$ that assigns a numerical score to each response. The preference between two responses is then determined solely by the reward scores for the two responses. For example, the Bradley-Terry (BT) model (Bradley & Terry, 1952) is a widely used method for modeling pairwise preferences in this context. However, the BT model can not capture intransitive (e.g. cyclic) preferences effectively (Bertrand et al., 2023). Under the BT model, the probability that response $\mathbf{y}$ is preferred over $\mathbf{y}^{\prime}$ is given by:
@@
$\max_{\theta}\ \mathbb{E}_{\mathbf{x}\sim\mathcal{X},\ \mathbf{y}\sim\pi_{\theta}(\cdot\mid\mathbf{x})}\left[r(\mathbf{y};\mathbf{x})\right]-\beta\mathbb{E}_{\mathbf{x}\sim\mathcal{X}}\left[\mathrm{KL}\left(\pi_{\theta}(\cdot\mid\mathbf{x})\,\|\,\pi_{\text{ref}}(\cdot\mid\mathbf{x})\right)\right],$ (3.1)
@@ 3.2 General Preference Modeling
$s(\mathbf{y}_{i}\succ\mathbf{y}_{j}\mid\mathbf{x}):=\log\frac{\mathbb{P}(\mathbf{y}_{i}\succ\mathbf{y}_{j}\mid\mathbf{x})}{1-\mathbb{P}(\mathbf{y}_{i}\succ\mathbf{y}_{j}\mid\mathbf{x})},$ (3.2)
@@ 3.2.1 Supervised Pair Preference Models
However, due to the language model’s position encoding (Press et al., 2021; Su et al., 2024) and the causal attention (Radford et al., 2018, 2019) mechanism not being symmetric, the candidate’s order in the concatenation will affect the final prediction results. It is mitigated by randomly shuffling the two responses in the training dataset but the output is still highly asymmetric. Another limitation is that how to represent the preference score can be highly ad-hoc. The two examples above already use different templates and different linear heads (sequential classification v.s. language modeling).
4 General Preference Modeling with Preference Representations
In this section, we propose a general preference representation learning framework that can model human preferences efficiently and expressively. Each response is embedded as a vector in a latent space, and the preferences are modeled through interactions between these representations (embeddings) using a skew-symmetric operator. We first define preference representations, which serve as the foundation for modeling the relationships between responses.
Definition 4.1 (Preference Representations).
Given a prompt $\mathbf{x}$, we assign to each response $\mathbf{y}$ a preference representation vector $\mathbf{v}_{\mathbf{y}|\mathbf{x}}\in\mathbb{R}^{2k}$. These representations are designed to capture the features relevant to human preferences beyond what can be represented by scalar rewards.
@@ 4 General Preference Modeling with Preference Representations
$\mathbf{R}_{l}=\begin{bmatrix}0&-1\\ 1&0\\ \end{bmatrix},\quad l=1,\dots,k.$ (4.1)
@@
Finally, we define the preference score, which quantifies the degree to which one response is preferred over another. This score is calculated based on the interaction between the preference representations, mediated by the skew-symmetric operator.
@@
The preference score between two responses $\mathbf{y}_{i}$ and $\mathbf{y}_{j}$ using preference representations is defined as:
@@
We model the preference probability using the logistic function as defined in Equation (3.3). Our general preference representation model (GPM) exhibits two desirable properties:
@@ 4.1 Expressiveness of the Model
Our general preference representation model is fully expressive for any real skew-symmetric preference matrix (see Appendix A.1 for complex representations interpretation). Specifically, we establish the following theorem (similar results have been proved in Balduzzi et al. (2018)):
Theorem 4.4 (Expressiveness of Preference Representation Model).
@@
Moreover, the vectors $\{\mathbf{v}_{i}\}$ can be constructed explicitly from $\mathbf{P}$.
Theorem 4.4 suggests that our preference representation framework can theoretically model arbitrary complex and potentially intransitive (e.g., cyclic) preference structures (see Appendix A.4 for proofs).
5 Efficient Preference Optimization with General Preference
$\max_{\pi}\ \mathbb{E}_{\mathbf{x}\sim\mathcal{X}}\left[\mathbb{E}_{\mathbf{y}\sim\pi(\cdot\mid\mathbf{x}),\ \mathbf{y}^{\prime}\sim\mu(\cdot\mid\mathbf{x})}\left[\mathbb{P}\left(\mathbf{y}\succ\mathbf{y}^{\prime}\mid\mathbf{x}\right)\right]-\beta\mathrm{KL}(\pi\|\pi_{\text{ref}})\right],$ (5.1)
$\max_{\pi}\ \min_{\pi^{\prime}}\ \mathbb{E}_{\mathbf{x}\sim\mathcal{X}}\left[\mathbb{E}_{\mathbf{y}\sim\pi(\cdot\mid\mathbf{x}),\ \mathbf{y}^{\prime}\sim\pi^{\prime}(\cdot\mid\mathbf{x})}\left[\mathbb{P}\left(\mathbf{y}\succ\mathbf{y}^{\prime}\mid\mathbf{x}\right)\right]\right].$ (5.2)
$\mathbb{P}\left(\pi\succ\pi^{\prime}\mid\mathbf{x}\right)=\mathbb{E}_{\mathbf{y}\sim\pi(\cdot\mid\mathbf{x}),\ \mathbf{y}^{\prime}\sim\pi^{\prime}(\cdot\mid\mathbf{x})}\left[\mathbb{P}\left(\mathbf{y}\succ\mathbf{y}^{\prime}\mid\mathbf{x}\right)\right].$ (5.3)
The optimization problem then becomes:
The von Neumann winner is a concept drawn from social choice theory (Sen, 1986) and has been studied in preference-based RL (Owen, 2013; Dudík et al., 2015). It is the Nash equilibrium of the two-player symmetric game (Equation 5.4). It represents a mixed strategy—a probability distribution over possible responses—that performs optimally in the worst-case scenario against any opponent.
More formally speaking, a distribution $\pi^{*}$ is called a von Neumann winner if it satisfies:
$\min_{\pi^{\prime}\in\Delta}\ \mathbb{E}_{\mathbf{x}\sim\mathcal{X}}\left[\mathbb{P}\left(\pi^{*}\succ\pi^{\prime}\mid\mathbf{x}\right)\right]\geq 1/2.$
This condition ensures that, on average, the von Neumann winner $\pi^{*}$ is at least as likely to be preferred than any other policy $\pi^{\prime}$. The von Neumann winner always exists due to the symmetric nature of the two-player game (Equation 5.4).
5.1 Efficient Policy Optimization
To align language models with the general preference, previous works (Wu et al., 2024b; Rosset et al., 2024; Liu et al., 2024; Swamy et al., 2024) have proposed iterative training frameworks that include 1) sampling from the current LLMs multiple responses $\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{K}$; 2) using a general preference model to label the preference among the $K$ responses; 3) aggregating the preferences into different learning objectives.
A key advantage of our model is its computational efficiency. The previous general preference models require $\mathcal{O}(K^{2})$ inference-time compute to evaluate all pairwise preferences among $K$ responses, as each pair is represented by a different concatenated sequence from $(\mathbf{x},\mathbf{y}_{i},\mathbf{y}_{j})$ to predict $\mathbb{P}(\mathbf{y}_{i}\succ\mathbf{y}_{j}|\mathbf{x})$. In contrast, computing the preference representation for $K$ responses requires only $O(K)$ forward passes / inference-time compute: we first calculate the representation $\mathbf{v}_{i}$ for each $\mathbf{y}_{i}$, and then use them to calculate the preference probability between any two responses using formula $s(\mathbf{y}_{i}\succ\mathbf{y}_{j})=\langle\mathbf{R}^{\succ}\mathbf{v}_{i},\mathbf{v}_{j}\rangle$. This way, our model is as efficient as a reward model while being way more expressive.
5.2 General Preference Optimization (GPO)

### v2

# Beyond Bradley-Terry Models: A General Preference Model for Language Model Alignment

@@ Abstract
Modeling human preferences is crucial for aligning foundation models with human values. Traditional reward modeling methods, such as the Bradley-Terry (BT) reward model, fall short in expressiveness, particularly in addressing intransitive preferences. In this paper, we introduce preference embedding, an approach that embeds responses into a latent space to capture intricate preference structures efficiently, achieving linear query complexity. Additionally, we propose preference score-based General Preference Optimization (GPO), which generalizes reward-based reinforcement learning from human feedback (RLHF). Experimental results show that our General Preference embedding Model (GPM) consistently outperforms the BT reward model on the RewardBench benchmark and effectively models cyclic preferences where any BT reward model behaves like a random guess. Furthermore, evaluations on downstream tasks such as AlpacaEval2.0, following the language model post-training with GPO and our general preference model, reveal performance improvements over BT models. These findings indicate that our method may enhance the alignment of foundation models with nuanced human values. The code is available at https://github.com/general-preference/general-preference-model.
@@ 1 Introduction
Modeling human preferences is a cornerstone in developing foundation models that interact seamlessly with users. In natural language modeling and reinforcement learning, aligning models with human intent and values has led to significant advancements, including improved text generation and enhanced decision-making policies (Ouyang et al., 2022; Christiano et al., 2017). Traditional approaches often rely on reward modeling, wherein a reward function is learned to guide the optimization of policies. While effective in certain contexts, these methods face expressiveness and computational efficiency challenges, particularly when addressing complex or intransitive human preferences (Tversky, 1969; Munos et al., 2023).
@@
(b) PairRM / PairPM
(c) General Preference embedding model (GPM)
Figure 1: Illustration of (a) Bradley Terry (BT) reward model, (b) supervised pair preference model (PairRM, PairPM) (Jiang et al., 2023; Dong et al., 2024), and (c) our General Preference embedding Model (GPM).
@@
On the other hand, supervised (sequential-classification) pair preference models (PairRM / PairPM) (Jiang et al., 2023; Dong et al., 2024) that predict the preference given a concatenation of the two responses can express complex and intransitive (cyclic) structures. But to fully capture the preference relations among $K$ responses, it requires evaluating $\mathcal{O}(K^{2})$ pairwise preferences between all $K$ candidate responses (Munos et al., 2023; Wu et al., 2024b). This quadratic scaling hinders them for applications with larger response sets especially in test-time scaling for reasoning tasks using verifiers and ranking models (Snell et al., 2024; Wu et al., 2024a).
In addition to computational inefficiency, supervised preference models exhibit asymmetric preference behaviors related to positions. The model’s design choice can also be highly ad hoc, varying among different templates and model architecture designs.
@@
In this paper, we answer this question affirmatively by proposing preference embedding, which bridges the gap between expressiveness and efficiency in general preference modeling. Our method embeds responses into a multi-dimensional latent space that captures the complex preference structure beyond transitive relations while allowing for efficient querying of preferences. Notably, our approach achieves a computational complexity of $\mathcal{O}(K)$, matching the efficiency of the BT model but with enhanced expressiveness.
@@
  • We introduce preference embedding for general preference modeling, enabling both efficient and expressive representation of human preferences. Our approach generalizes the Bradley-Terry (BT) reward model by embedding responses into a latent space, capturing complex structures, including intransitive preferences. Notably, our General Preference embedding model (GPM) achieves a query complexity of $\mathcal{O}(K)$ for evaluating preferences among $K$ responses which match the complexity of the Bradley-Terry reward model, an improvement over the $\mathcal{O}(K^{2})$ complexity of traditional supervised preference models that rely on pairwise inputs (see Section 4).
  • We demonstrate GPM’s effectiveness across various tasks, including CyclicPreference (ours) and the renowned RewardBench (Lambert et al., 2024). Specifically, GPM models intransitive (e.g., cyclic) preferences with near-perfect accuracy, whereas the BT reward model performs like random guessing (see Section 6.2). Additionally, GPM consistently outperforms the BT reward model on RewardBench (see Section 6.1).
  • For language model alignment, we propose General Preference Optimization (GPO), which leverages the preference scores provided by GPM. The general preference score can also be integrated as a preference signal into a wide range of RLHF and preference optimization methods (Rafailov et al., 2024; Munos et al., 2023; Wu et al., 2024b). Experimental results on AlpacaEval-2.0 reveal that our approach may improve reward-based language model alignment methods (see Section 6.3).
@@ 2 Related Work
Figure 2: Intransitiveness in real-world preferences.
Reward-Based Reinforcement Learning from Human Feedback (RLHF).
Typical approaches to modeling human preference for language model alignment usually learn a reward model from a preference dataset. The human preference is assumed to follow the Bradley-Terry (BT) model (Bradley & Terry, 1952) or the Thurstone model (Thurstone, 2017). LLM policies then are fine-tuned to maximize these scalar reward signals for better alignment (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022). Later, the direct preference optimization (DPO) methods are proposed by Rafailov et al. (2024) only implicitly to learn a reward model represented by an LLM. The human preference is still assumed to follow the Bradley-Terry model. However, the reliance on scalar rewards imposes a total ordering on preferences, which may not reflect the intransitive or stochastic nature of human judgments (Tversky, 1969; Agranov & Ortoleva, 2017).
@@ 3 Background
In this section, we present preliminaries on reward modeling, preference modeling, and reinforcement learning from human feedback (RLHF) for language model alignment. We consider an autoregressive language model that generates responses to the given prompts. Let $\mathbf{x}=[x_{1},x_{2},\ldots]$ denote a prompt (a sequence of tokens). The language model $\pi$ generates a response $\mathbf{y}=[y_{1},y_{2},\ldots,y_{N}]$ based on the conditional probability distribution: $\pi(\mathbf{y}\mid\mathbf{x})=\prod_{i=1}^{N}\pi\left(y_{i}\mid\mathbf{x},\mathbf{y}_{<i}\right)$, where $\mathbf{y}_{<i}$ represents the sequence of tokens generated before position $i$. In this paper, we assume a general-preference oracle. Given two responses $\mathbf{y}$ and $\mathbf{y}^{\prime}$ to the same prompt $\mathbf{x}$, the oracle provides the feedback indicating which response is preferred.
@@ 3.1 Reward-Based Reinforcement Learning from Human Feedback
The most prevalent approach to aligning language models with human preferences is to consider a scalar reward function $r(\mathbf{y};\mathbf{x})$ that assigns a numerical score to each response. The preference between two responses is then determined solely by the reward scores for the two responses. For example, the Bradley-Terry (BT) model (Bradley & Terry, 1952) is a widely used method for modeling pairwise preferences in this context. However, the BT model can not capture intransitive (e.g. cyclic) preferences effectively (Bertrand et al., 2023). Under BT model, the probability that response $\mathbf{y}$ is preferred over $\mathbf{y}^{\prime}$ is given by:
@@
$\max_{\theta}\ \mathbb{E}_{\mathbf{x}\sim\mathcal{X},\ \mathbf{y}\sim\pi_{\theta}(\cdot\mid\mathbf{x})}$ $\left[r(\mathbf{y};\mathbf{x})\right]-$
$\beta\mathbb{E}_{\mathbf{x}\sim\mathcal{X}}\left[\mathrm{KL}\left(\pi_{\theta}(\cdot\mid\mathbf{x})\,\|\,\pi_{\text{ref}}(\cdot\mid\mathbf{x})\right)\right],$
@@ 3.2 Preference Modeling
$s(\mathbf{y}_{i}\succ\mathbf{y}_{j}\mid\mathbf{x}):=\log\frac{\mathbb{P}(\mathbf{y}_{i}\succ\mathbf{y}_{j}\mid\mathbf{x})}{1-\mathbb{P}(\mathbf{y}_{i}\succ\mathbf{y}_{j}\mid\mathbf{x})},$
@@ 3.2.1 Supervised Pair Preference Models
However, due to the language model’s position encoding (Press et al., 2021; Su et al., 2024) and the causal attention (Radford et al., 2018, 2019) mechanism not being symmetric, the candidate’s order in the concatenation will affect the final prediction results. It is mitigated by randomly shuffling the two responses in the training dataset but the output is still highly asymmetric. Another limitation is that how to represent the preference score can be highly ad-hoc. The two examples above already use different templates and different model architectures (sequential classification v.s. language modeling).
3.3 Preference-based Reinforcement Learning from Human Feedback
$\max_{\pi}\ \mathbb{E}_{\mathbf{x}\sim\mathcal{X}}\bigg{[}$ $\mathbb{E}_{\mathbf{y}\sim\pi(\cdot\mid\mathbf{x}),\ \mathbf{y}^{\prime}\sim\mu(\cdot\mid\mathbf{x})}\Big{[}\mathbb{P}\big{(}\mathbf{y}\succ\mathbf{y}^{\prime}\mid\mathbf{x}\big{)}\Big{]}-\beta\,\mathrm{KL}\big{(}\pi\|\pi_{\text{ref}}\big{)}\bigg{]},$
$\max_{\pi}\ \min_{\pi^{\prime}}\ \mathbb{E}_{\mathbf{x}\sim\mathcal{X}}f\left[\mathbb{E}_{\mathbf{y}\sim\pi(\cdot\mid\mathbf{x}),\ \mathbf{y}^{\prime}\sim\pi^{\prime}(\cdot\mid\mathbf{x})}\left[\mathbb{P}\left(\mathbf{y}\succ\mathbf{y}^{\prime}\mid\mathbf{x}\right)\right]\right].$
$\mathbb{P}\left(\pi\succ\pi^{\prime}\mid\mathbf{x}\right)=\mathbb{E}_{\mathbf{y}\sim\pi(\cdot\mid\mathbf{x}),\ \mathbf{y}^{\prime}\sim\pi^{\prime}(\cdot\mid\mathbf{x})}\left[\mathbb{P}\left(\mathbf{y}\succ\mathbf{y}^{\prime}\mid\mathbf{x}\right)\right].$
The optimization problem then becomes:
4 General Preference Embedding Model
In this section, we propose a general preference embedding framework that can efficiently and expressively model human preferences. Each response is embedded as a vector in a latent space, and the preferences are modeled through interactions between these embeddings using a skew-symmetric operator. We first define preference embeddings, which serve as the foundation for modeling the relationships between responses.
Definition 4.1 (Preference Embeddings).
Given a prompt $\mathbf{x}$, we assign to each response $\mathbf{y}$ a preference embedding vector $\mathbf{v}_{\mathbf{y}|\mathbf{x}}\in\mathbb{R}^{2k}$. These embeddings are designed to capture the features relevant to human preferences beyond what can be represented by scalar rewards.
@@ 4 General Preference Embedding Model
$\mathbf{R}_{l}=\begin{bmatrix}0&-1\\ 1&0\\ \end{bmatrix},\quad l=1,\dots,k.$
@@
Finally, we define the preference score, which quantifies the degree to which one response is preferred over another. This score is calculated based on the interaction between the preference embeddings, mediated by the skew-symmetric operator.
@@
The preference score between two responses $\mathbf{y}_{i}$ and $\mathbf{y}_{j}$ using preference embeddings is defined as:
@@
We model the preference probability using the logistic function as defined in (3.1)). Our general preference embedding model (GPM) exhibits two desirable properties:
@@ 4.1 Expressiveness of the Model
Our general preference embedding model is fully expressive for any real skew-symmetric preference matrix (see Appendix A.1 for complex representations interpretation). Specifically, we establish the following theorem (similar results have been proved in Balduzzi et al. (2018)):
Theorem 4.4 (Expressiveness of Preference Embedding Model).
@@
Theorem 4.4 suggests that our preference embedding framework can theoretically model arbitrary complex and potentially intransitive (e.g., cyclic) preference structures (see Appendix C for proofs).
4.2 Implementing General Preference Embedding Model
Theorem 4.5.
Let $\mathbf{P}\in\mathbb{R}^{2k\times 2k}$ be a real skew-symmetric matrix (i.e., $\mathbf{P}=-\mathbf{P}^{\top}$). Then there exist embeddings $\{\mathbf{v}_{i}\}_{i=1}^{2k}\subset\mathbb{R}^{2k}$ and a block-diagonal skew-symmetric matrix $\mathbf{R}^{\succ}\in\mathbb{R}^{2k\times 2k}$, such that:
To effectively capture general preferences while maintaining computational efficiency, we implement our preference embedding model by augmenting an existing language model with two additional components: an eigenvalue scale gate and an eigenvector embedding head.
$\{\lambda_{l}(\mathbf{x})\}=\mathcal{G}_{\lambda}(\mathbf{x}).$
where $\mathbf{D}(\mathbf{x})$ is a block-diagonal matrix with blocks $\sqrt{\lambda_{l}(\mathbf{x})}\mathbf{I}_{2}$, and $\mathbf{R}^{\succ}$ is the skew-symmetric preference operator. We normalize the embeddings $\mathbf{v}_{\mathbf{y}}$ to have unit length to ensure training stability.
5 General Preference Optimization